A crate of eggs is located in the middle of the flat bed of a pickup truck as the truck negotiates an unbanked curve in the road. The curve may be regarded as an arc of a circle of radius r = 35.0m. If the coefficient of static friction between crate and truck is 0.600, how fast can the truck be moving without the crate sliding?


1
Expert's answer
2021-05-31T15:38:12-0400
"\frac{mv^2}{r}=\mu mg\\v=\sqrt{\mu r g}\\v=\sqrt{(0.6)(35)(9.8)}=14.3\frac{m}{s^2}"